Lots more changes have gone in to this latest version since the last testing at the end of December. Mostly, I wanted to check out moving the seeing calculation to every 5 seconds - to make sure that didn't break anything. The second priority was to test the offsetting that didn't work on-sky last test (although the test was sketchy last time, we don't know what the telescope was doing at the time we saw weird behavior.)
They didn't open the shutters until later than normal because of MCSPU problems.
Juan Carlos was happy with the adjusted seeing and R0 calculations in the log.
Had to work on getting FITS files out when I wanted them. finished around 9pm
Spiral testing still showed the same problems (of course I didn't put any code in to fix it, so makes sense)
Started offset testing about 9:22.
DEC offset wasn't calculated correctly, always 0
resetting base still too often - had to modify code to only set base on the slew to a star
about 20-25 minutes just trying to get DIMM to acquire - cannot test offset unless you actually find the star
the RA offset seemed to be applied correctly, but we always ended up with the star above the target on the web cam - even when I changed the sign on the application of the DEC offset
the offset seems to accumulate, getting worse in DEC